# VLMQ: Efficient Post-Training Quantization for Large Vision-Language Models via Hessian Augmentation

### v1

@@ 5 VLMQ
To address the importance discrepancy across tokens, we propose an accurate PTQ method for VLMs by optimizing an importance-aware objective, dubbed VLMQ. Specifically, 1) to capture the varying salience across tokens, we refine the quantization objective to an importance-aware one, yielding an enhanced Hessian. Then, 2) we establish a theoretical link between block-wise loss perturbation and layer-wise output errors, enabling efficient computation of gradient-driven importance scores via a single block-wise backward pass.
Importance-Aware Objective
Although Equation 4 yields the optimal quantized weight $\hat{{\mathbf{w}}}$ by minimizing the squared error, it inherently assumes equal contribution across all tokens in ${\mathbf{Z}}={\mathbf{W}}{\mathbf{X}}$. Such equal treatment may introduce bias due to the vision over-representation. To address this limitation, we introduce a token-level importance weighting matrix ${\mathbf{G}}\in{\mathbb{R}}^{N\times N}$, where $\mathbf{G}$} is diagonal with the $i$-th diagonal element representing the importance assigned to output token ${\mathbf{Z}}_{:,i}$. Incorporating importance factors into the objective leads to a refined formulation:
@@
which gives the corresponding optimal weight updating as
$\Delta{\mathbf{w}}$ $=\frac{\left(\hat{{\mathbf{w}}}_{q}-{\mathbf{w}}_{q}\right)}{\left[{\mathbf{X}}{\mathbf{G}}{\mathbf{X}}^{\top}\right]_{qq}^{-1}}\cdot\widetilde{{\mathbf{H}}}_{q:,}^{-1}+{\mathbf{r}}{\mathbf{G}}{\mathbf{X}}^{\top}\left[{\mathbf{X}}{\mathbf{G}}{\mathbf{X}}^{\top}\right]_{-q}^{-1}$ (7)
where $\widetilde{{\mathbf{H}}}={\mathbf{X}}{\mathbf{G}}{\mathbf{X}}^{\top}$, $\tilde{{\mathbf{r}}}={\mathbf{r}}\sqrt{\mathbf{G}}$ and $\widetilde{{\mathbf{X}}}={\mathbf{X}}\sqrt{\mathbf{G}}$ indicate the importance-aware Hessian, residual, and activation matrix yielded from the refined objective (Equation 5). Our importance-aware weight update presented in Equation 7 remains formally consistent with the original GPTAQ (Li et al. 2025b) formulation. Distinct from BoA (Kim et al. 2024), which attributes importance at the channel level, our approach assigns importance scores at the token-level activation granularity. This design is orthogonal to the parallelization strategies proposed in GPTQ (Frantar et al. 2022) and GPTAQ (Li et al. 2025b). The compatibility allows us to directly inherit efficiency tricks (e.g., Cholesky reformulation) proposed therein.
Gradient-Driven Importance Factors
Figure 4: Visualization of importance factors of k_proj and v_proj outputs.
Figure 3 illustrates the overall process of computing gradient-driven importance factors, which involves three key stages: Forward, Backward, and Gradient Processing. We begin by introducing the forward pass and defining the block-wise target. Next, we theoretically justify the use of gradients for importance estimation and describe how gradient caching is performed during backpropagation. Finally, we present the formal procedure for computing token-level importance factors from the raw gradients.
Forward. During Forward, we place activation hooks immediately after attention modules to extract hidden states. This enables the computation of a localized loss ${\mathcal{L}}_{\text{Block}}$ between the semi-quantized model and its auxiliary full precision counterpart, which is used to trigger a one-time localized backward pass per block. Here, to balance efficiency and accuracy in deriving importance scores, we define a block as an attention module as formulated in Equation 1, which encompasses multi-head self-attention along with Q/K/V/O projections and the residual stream. Additionally, we employ the block output mean square error (MSE) as the localized objective as
@@
where $\mathbf{X}$} and $\hat{{\mathbf{X}}}$ are the full precision activations and activations with previous layers quantized, respectively. While it is theoretically possible to treat a full decoding layer as a block, doing so often incurs substantial memory overhead (Ding et al. 2023), rendering the method impractical on resource-constrained hardware.
Backward. The Backward module illustrates the internal mechanics within an attention layer. During this step, we perform a single pass of backpropagation on ${\mathcal{L}}_{\text{Block}}$ defined by Equation 8 and cache the gradients of each linear projection output (i.e., Q/K/V/O projections) to estimate token-level importance factors. We propose to use gradient-driven importance factors to quantify the contribution of each token, which is supported by Theorem 5.1.
Theorem 5.1.
The block-wise loss perturbation $\Delta{\mathcal{L}}_{\text{Block}}$ can be approximate by the first-order error as
$\Delta{\mathcal{L}}_{\text{Block}}\approx\Delta\theta{\mathbf{p}}^{(\Delta\theta),\top}+{\mathcal{O}}(|\theta|^{2})\approx\Delta{\mathbf{z}}{\mathbf{p}}^{(\Delta{\mathbf{z}}),\top},$ (9)
The proof is provided in Appendix A. To this end, the gradient ${\mathbf{p}}^{(\Delta{\mathbf{z}})}$ measures the disturbance of layer output error on block loss, which can be an effective metric for token importance. Henceforth, we omit the superscript of ${\mathbf{p}}^{(\Delta{\mathbf{z}})}$ and reshape it to ${\mathbf{P}}\in{\mathbb{R}}^{C_{o}\times N}$ as our raw gradients.
Process Grad. Finally, the Process Grad module illustrates how we construct the importance factor matrix ${\mathbf{G}}\in{\mathbb{R}}^{N\times N}$ from the raw gradients $\mathbf{P}$}. Formally, the importance factor is defined as:
${\mathbf{G}}=\operatorname{Diag}\left(\left[\overline{\left|{\mathbf{P}}\right|}_{0},\,\overline{\left|{\mathbf{P}}\right|}_{1},\,\dots,\,\overline{\left|{\mathbf{P}}\right|}_{N-1}\right]\right),$ (10)
where the importance of $n$-th token is the $\ell_{1}$-norm of $n$-th column in gradients:
$\overline{\left|{\mathbf{P}}\right|}_{n}=\frac{1}{C_{o}}\sum_{i=0}^{C_{o}-1}\left|{\mathbf{P}}\right|_{i,n}.$ (11)
Note that importance factors are also computed for text tokens to preserve the relative contribution balance across modalities. These gradient-driven importance factors not only reflect inter-layer dependencies but also implicitly indicate how to adjust for cumulative errors introduced by previous quantized layers. Figure 4 visualizes importance factors of k_proj and v_proj output within a representative attention block. It is evident that the distribution of token salience exhibits distinct patterns across layers. For instance, several vision tokens are identified as salient in the k_proj output, whereas in the v_proj output, vision tokens are predominantly assigned low importance, indicating redundancy. This distribution highlights the diverse contribution of vision tokens to block-wise loss and further echoes the necessity of applying fine-grained, importance-aware calibration. To quantize the full model, we progressively alternate between local backpropagation and calibration, which better captures error propagation dynamics. The VLMQ is outlined in Appendix C.
6 Experiments
@@
Qwen2-VL-2B-Instruct-INT3
Full Prec 72.61 89.35 40.27 43.59 76.60 74.23 61.79 79.38 67.23
GPTQ 58.84 74.55 33.25 30.83 61.50 65.13 50.37 73.49 56.00
GPTAQ 58.76 74.03 38.71 28.71 61.10 64.25 52.17 73.18 56.36
@@
Qwen2-VL-7B-Instruct-INT3
Full Prec 81.44 93.89 57.30 56.09 80.90 85.55 69.26 82.02 75.81
GPTQ 72.80 88.35 47.76 27.90 72.40 78.26 63.02 78.28 66.10
GPTAQ 72.76 88.09 49.76 43.96 73.80 80.41 64.21 79.23 69.03
[…]
LLaVA-OneVision-7B-INT3
Full Prec 80.08 87.09 57.39 53.93 62.10 89.98 64.82 75.96 71.42
GPTQ 74.24 77.82 46.47 42.59 56.60 85.26 60.65 71.90 64.44
GPTAQ 75.72 78.07 46.48 42.94 58.30 85.97 61.84 72.69 65.25
[…]
Table 2: INT3 performance comparison across different quantization methods on eight benchmarks, which are ChartQA, DocVQA (Validation set), MME-RealWorld (English), MME-RealWorld (Chinese), OCRBench, ScienceQA, SeedBench 2 Plus, and TextVQA (Validation set), respectively. Best and second-best results are highlighted in bold and underline formats, respectively. indicates that we use the token-level $\ell_{2}$-norm of row gradients instead of $\ell_{1}$-norm as our importance factors.
[…]
Qwen2.5-VL-7B-Instruct-INT2g128
GPTQ 57.72 78.79 44.94 13.89 70.80 55.77 48.40 74.36 55.58
GPTAQ 53.48 74.82 39.38 3.60 67.70 2.95 17.83 73.79 41.69
@@
Implementation Details
Models. We conduct experiments on open-source SOTA VLMs, including LLaVA-OneVision-0.5B/7B, Qwen2.5-VL-2B/7B/32B-Instruct, and Qwen2-VL-7B-Instruct. For the LLaVA-OneVision series, we select versions adopting Qwen2 as LM backbones and SigLIP-400M (Zhai et al. 2023) as their vision encoder.
Calibration. Following MBQ (Li et al. 2025a), we select the improved COCO Caption dataset introduced by ShareGPT4V (Chen et al. 2024b). We randomly sample 512 text-image pairs as our calibration set. Unless otherwise specified, we use the default dampening ratio $\lambda=0.01$ and enable act_order for all experiments. Other details can be found in Appendix D.
Evaluation. We undertake an extensive evaluation of quantized VLMs using multiple challenging vision-language tasks based on the LMMs-Eval framework (Zhang et al. 2024a). Specifically, we employ representative tasks including ChartQA (Masry et al. 2022), DocVQA (Mathew, Karatzas, and Jawahar 2021), MME-RealWorld (Zhang et al. 2024b), OCRBench (Liu et al. 2024), ScienceQA (Lu et al. 2022), SeedBench 2 Plus (Li et al. 2024a), and TextVQA (Singh et al. 2019), thereby comprehensively measuring the text recognition, visual perception, and visual reasoning capabilities of quantized models. Flash Attention (Dao et al. 2022) is enabled, and default configurations are used throughout all experiments.
Main Results
INT3 quantization. We assess the zero-shot performance of INT3 quantized models across eight vision-language benchmarks in Table 2. VLMQ demonstrates overall superiority compared to baseline approaches and achieves up to 1.01% average accuracy improvement. For Qwen2.5-VL-7B-Instruct-INT3, the proposed VLMQ outperforms its counterparts quantized by GPTQ and GPTAQ on all benchmarks except MME-RealWorld (English). Remarkably, it yields accuracy improvements of 1.64%, 1.40%, and 1.18% on ChartQA, MME-RealWorld (Chinese), and ScienceQA, respectively. For the LLaVA-OneVision-7B-INT3 model, we apply the token-level $\ell_{2}$-norm instead of the $\ell_{1}$-norm adopted in the other experiments, which we empirically find beneficial for performance enhancement.
INT2g128 quantization. We further evaluate the performance of INT2 group-wise quantized models with group_size=128 as shown in Table 3. While taking GPTAQ as our precursor algorithm generally yields satisfactory results under INT3 quantization, applying this configuration to INT2g128 leads to severe performance degradation. In particular, GPTAQ produces even worse results than GPTQ under this ultra-low-bit setting. Taking Qwen2-VL-7B-Instruct-INT2g128 as an example, GPTAQ achieves only 41.69% accuracy, significantly lower than the 55.58% achieved by GPTQ. Additionally, the Qwen2.5-VL-7B-Instruct-INT2g128 fails on MME-RealWorld (Chinese) and ScienceQA benchmarks, achieving only 3.60% and 2.95% accuracy, respectively. A plausible explanation is that, under extremely low-bit quantization, the residual error in GPTAQ propagated from previously quantized layers may be continuously accumulated and significantly amplified. This accumulated noise severely interferes with the dominant weight update term $\left(\hat{{\mathbf{w}}}_{q}-{\mathbf{w}}_{q}\right)/{\mathbf{H}}_{qq}^{-1}$, thereby compromising the undesirable quantization quality. To mitigate this issue, we choose GPTQ as our foundation algorithm. Evidently, we observe a notable performance improvement of VLMQ under the INT2g128 configuration. For instance, the MME-RealWorld (Chinese) benchmark exhibits a notable 16.45% accuracy gain for Qwen2.5-VL-7B-Instruct-INT2g128 compared with GPTQ.
Ablation Study
Bit Calib Data Bias Layer Acc
LLaVA-OneVision-7B
FP - - 71.42
INT3 WikiText None 65.01
COCO None 65.25
COCO Q/K/V Proj 64.73
COCO Q/K/V/O Proj 65.69
Qwen2-VL-7B-Instruct-INT3
FP - - 75.81
INT3 WikiText None 66.93
COCO None 69.03
COCO Q/K/V Proj 56.21
COCO Q/K/V/O Proj 69.42
Table 4: Ablation studies on LLaVA-OneVision-7B and Qwen2-VL-7B-Instruct.
We conduct step-by-step ablation studies on Qwen2-VL-7B-Instruction-INT3 and LLaVA-OneVision-INT3, as summarized in Table 4. It is evident that using WikiText alone as the calibration dataset leads to substantial performance degradation, underscoring the necessity of involving vision modality at quantization time. By incorporating token-level importance variance, VLMQ augments the Hessian matrix with importance-aware weighting, thereby improving the weight update step $\Delta{\mathbf{W}}$ and enhancing overall quantization quality. Notably, we find that applying our biased calibration strategy to o_proj is critical. We attribute this to the error accumulation caused by quantizing q_proj, k_proj, and v_proj. Leveraging enhanced Hessian information in o_proj allows effective error compensation within each residual block, ultimately improving quantization fidelity.
Quantization Efficiency
Model Size GPU Hours

### v2

@@ 4 VLMQ
Figure 3: Visualization of normalized token-wise error ($\Delta{\mathbf{z}}$) and gradient (${\mathbf{p}}^{(\Delta{\mathbf{z}})}$). The red circle indicates the salient vision tokens. The magnitude of the error remains relatively stable across tokens, whereas the gradient varies across tokens in different modalities.
Considering the aforementioned limitations, we propose VLMQ, an accurate PTQ framework for VLMs that calibrates quantization in an importance-aware manner. Concretely, 1) to identify salient tokens, we introduce a diagonal gradient-derived importance factor, where each element reflects token-level importance. Then, 2) to compute this factor efficiently, we acquire raw gradients through a single lightweight block-wise backpropagation. Furthermore, 3) we then refine the optimization objective by incorporating the proposed importance factor, enabling selective emphasis on salient tokens during quantization.
4.1 Effective Gradient-Driven Importance Factor
To quantify the contribution of tokens across modalities during calibration, we introduce our gradient-driven and token-level importance factors as below.
Theorem 4.1
The target loss perturbation $\Delta{\mathcal{L}}$ can be approximated by the first-order error as
$\Delta{\mathcal{L}}\approx\Delta\theta{\mathbf{p}}^{(\Delta\theta),\top}+{\mathcal{O}}(|\theta|^{2})\approx\Delta{\mathbf{z}}{\mathbf{p}}^{(\Delta{\mathbf{z}}),\top},$ (3)
The proof is provided in Appendix 0.D. Theorem 4.1 demonstrates that the loss perturbation $\Delta{\mathcal{L}}$ is influenced by two contributors: the output errors ($\Delta{\mathbf{z}}$) and the gradients (${\mathbf{p}}^{(\Delta{\mathbf{z}})}$). A visualization of token-wise errors and gradients is shown in Figure 3. It is clear that the magnitudes of these two factors are not necessarily correlated. Specifically, although the errors attributed to different tokens are very similar, there is a clear disparity in their corresponding gradients. For instance, the gradients for redundant vision tokens are considerably smaller than those for important text tokens. Unlike previous studies [lin2023awq, frantar2022gptq] that optimize quantization only by minimizing layer-wise MSE (i.e. $\Delta{\mathbf{z}}||^{2}$), in this work, we incorporate this gradient information to capture the variance of importance between tokens. This helps separately consider different modalities at the token level while downscaling the importance of redundant vision tokens during quantization. For simplicity, we omit the superscript of ${\mathbf{p}}^{(\Delta{\mathbf{z}})}$ and denote it for a certain layer as ${\mathbf{P}}\in{\mathbb{R}}^{C_{o}\times N}$ in the following paragraphs.
4.1.1 Gradient processing.
We now describe the acquisition of the token-wise importance factor $\mathbf{G}$} (a diagonal matrix) from the raw gradients $\mathbf{P}$}. Formally, the importance factor is defined as:
${\mathbf{G}}=\operatorname{Diag}\left(\left[\overline{\left|{\mathbf{P}}\right|}_{0},\,\overline{\left|{\mathbf{P}}\right|}_{1},\,\dots,\,\overline{\left|{\mathbf{P}}\right|}_{N-1}\right]\right),\ \text{where}\ \overline{\left|{\mathbf{P}}\right|}_{n}=\frac{1}{C_{o}}\sum_{i=0}^{C_{o}-1}\left|{\mathbf{P}}\right|_{i,n}.$ (4)
Figure 4 illustrates this process. We convert the full raw gradients into a diagonal importance factor $\mathbf{G}$}. To validate the effectiveness of $\mathbf{G}$}, alternative importance factors, such as attention score-style metrics [chen2024fastv, dhouib2025pact], have been considered. However, they fail to accurately reflect token-wise importance variance across modalities, resulting in substantial performance degradation (see Section 5.3).
Figure 4: Derivation of Visualization of importance factors.
4.2 Efficient Gradient Acquisition
To extract raw gradients, we consider three potential types of target loss: i.e. 1) layer-wise distillation loss, 2) block-wise distillation loss, and 3) network-wise supervised fine-tuning (SFT) loss. The layer-wise approach is relatively efficient but fails to capture cross-layer dependencies. The network-wise approach provides a more accurate gradient information, yet it incurs prohibitive computational cost and risks overfitting the quantized model to the limited calibration dataset [gong2025pushing]. Details can be found in Section 5.3. To balance efficiency and effectiveness, we adopt a compromise solution by employing the block-wise approach.
4.2.1 Efficient block-wise backpropagation.
We place activation hooks immediately after attention modules to extract hidden states. This enables the computation of a localized loss ${\mathcal{L}}_{\text{Block}}$ between the semi-quantized model and its auxiliary full precision counterpart, which is used to trigger a one-time localized backward pass per block. Here, we define a block as an attention module as formulated in equation 1, which encompasses multi-head self-attention along with Q/K/V/O projections and the residual stream. Additionally, we employ the block MSE as the localized objective as
@@ 4.2 Efficient Gradient Acquisition
where $\mathbf{X}$} and $\hat{{\mathbf{X}}}$ denote the full-precision activations and the activations with preceding layers quantized, respectively. Although in theory one could define an entire decoding layer as a block, this choice typically incurs prohibitive memory overhead [ding2023cbq], making the approach impractical for resource-constrained hardware.
Figure 5: Pipeline of computing importance factors. The “Forward” module illustrates the quantization dataflow across decoding layers, where a breakpoint is set at the output of each attention module to compute the local loss ${\mathcal{L}}_{\text{Block}}$ and trigger a localized backward pass. The “Backward” module details the internal operations within an attention block, where gradients of each linear projection output are cached to derive token-level importance factors.
4.3 Importance-Aware Objective
We now present our refined optimization objective incorporating the importance factor $\mathbf{G}$}. In contrast to prior works [lin2023awq, frantar2022gptq] that minimize the output MSE, we adopt an importance-aware formulation, formally defined as
@@ 4.3 Importance-Aware Objective
Note that this formulation can be seamlessly integrated into diverse PTQ frameworks. In this work, we adopt GPTAQ [li2025gptqv2] as our precursor algorithm due to its simplicity and ease of implementation. To quantize the entire model with the proposed VLMQ, we progressively alternate between lightweight block-wise backpropagation and importance-aware calibration, thereby capturing error propagation dynamics more effectively. The detailed framework formulation and theoretical derivations are presented in Appendix 0.C.
5 Experiments
5.1 Implementation Details
5.1.1 Models.
We conduct experiments on open-source SOTA VLMs, including LLaVA-OneVision-0.5B/7B, Qwen2.5-VL-2B/7B/32B-Instruct, and Qwen2-VL-7B-Instruct. For the LLaVA-OneVision series, we select versions adopting Qwen2 as LM backbones and SigLIP-400M [zhai2023siglip] as their vision encoder.
5.1.2 Calibration.
We select the improved COCO Caption dataset introduced by ShareGPT4V [chen2024sharegpt4v]. We randomly sample 512 text-image pairs as our calibration set. Unless otherwise specified, we use the default dampening ratio $\lambda=0.01$ and enable act_order for all experiments. Other details can be found in Appendix 0.E.
5.1.3 Evaluation.
We undertake an extensive evaluation of quantized VLMs using multiple challenging vision-language tasks based on the LMMs-Eval framework [zhang2024lmms-eval]. Specifically, we employ representative tasks including ChartQA [masry2022chartqa], DocVQA [mathew2021docvqa], MME-RealWorld [zhang2024mme-realword], OCRBench [liu2024ocrbench], ScienceQA [lu2022scienceqa], SeedBench 2 Plus [li2024seed2plus], and TextVQA [singh2019textvqa], thereby comprehensively measuring the text recognition, visual perception, and visual reasoning capabilities of quantized models. Flash Attention [dao2022flashattention] is enabled, and default configurations are used throughout all experiments.
5.2 Main Results
5.2.1 INT3g128 quantization.
Table 2: INT3g128 performance comparison across different quantization methods on eight benchmarks, which are ChartQA, DocVQA (Validation set), MME-RealWorld (English), MME-RealWorld (Chinese), OCRBench, ScienceQA, SeedBench 2 Plus, and TextVQA (Validation set), respectively.
@@ 5.2 Main Results
Qwen2-VL-2B-Instruct-INT3g128
Full Prec 72.61 89.35 40.27 43.59 76.60 74.23 61.79 79.38 67.23
AWQ 62.96 83.96 30.56 31.89 66.90 65.95 53.67 73.23 58.64
MBQ 62.80 82.07 28.96 33.45 66.30 67.72 54.98 73.40 58.71
GPTQ 62.60 85.29 37.80 35.54 70.50 67.15 58.15 77.23 61.78
GPTAQ 63.60 84.92 44.12 42.10 71.60 66.61 56.79 77.81 63.44
VLMQ 67.00 85.06 37.75 42.40 69.30 67.20 56.92 77.55 62.90
Qwen2-VL-7B-Instruct-INT3g128
Full Prec 81.44 93.89 57.30 56.09 80.90 85.55 69.26 82.02 75.81
AWQ 77.60 93.46 56.72 49.97 77.40 82.39 66.84 78.91 72.91
MBQ 77.20 93.68 56.79 50.50 76.60 82.24 65.74 78.67 72.68
GPTQ 79.08 92.21 54.47 49.21 79.80 83.16 66.80 81.02 73.22
GPTAQ 78.92 92.29 55.83 52.12 79.00 83.05 67.28 80.98 73.68
@@
LLaVA-OneVision-7B-INT3g128
Full Prec 80.08 87.09 57.39 53.93 62.10 89.98 64.82 75.96 71.42
GPTQ 78.40 84.53 55.20 49.64 60.30 88.85 63.46 74.55 69.37
GPTAQ 78.64 84.06 54.82 50.53 59.80 88.26 64.82 74.76 69.46
@@
We evaluate the zero-shot performance of INT3g128 (group_size=128) quantized models across eight vision-language benchmarks, as shown in Table 2. VLMQ demonstrates its overall superiority across various benchmarks. Notably, for Qwen2-VL-7B-Instruct-INT3g128, the proposed VLMQ achieves accuracies of 57.34% and 55.47% on MME-RealWorld (English/Chinese), respectively, and maintains competitive performance across other QA tasks, showcasing its robustness and versatility. Furthermore, for other model variants, VLMQ continues to narrow the gap between full-precision models and their quantized counterparts. For Qwen2-VL-2B-Instruct-INT3g128, VLMQ generally achieves improved or comparable performance across various benchmarks. However, it slightly lags behind GPTAQ, particularly due to GPTAQ’s superior performance on MME-RealWorld (English), where GPTAQ outperforms the full-precision model by 3.85%. We also present the INT3 (w/o group-wise quantization) quantization comparison table in Appendix 0.E, where VLMQ demonstrates notable performance improvements across multiple benchmarks, further highlighting its effectiveness in various settings.
5.2.2 INT2g128 quantization.
@@
Qwen2.5-VL-7B-Instruct-INT2g128
GPTQ 57.72 78.79 44.94 13.89 70.80 55.77 48.40 74.36 55.58
GPTAQ 53.48 74.82 39.38 3.60 67.70 2.95 17.83 73.79 41.69
[…]
We further evaluate the performance of INT2 group-wise quantized models with group_size=128 as shown in Table 3. Under these ultra-low-bit settings, transformation-based methods including AWQ [lin2023awq] and MBQ [li2025mbq] fail to produce desirable results. We choose GPTQ as our foundation algorithm for some experiments (see detailed explanation in Appendix 0.E). Evidently, we observe a significant performance improvement of VLMQ under the INT2g128 configuration. For instance, the MME-RealWorld (Chinese) benchmark demonstrates a substantial 16.45% accuracy gain for Qwen2.5-VL-7B-Instruct-INT2g128 compared to GPTQ. This clear enhancement under the INT2g128 setting further validates the effectiveness of VLMQ in ultra-low-bit quantization, highlighting its ability to maintain model performance even with highly constrained bit-widths.
5.2.3 Generalized to more tasks.
Table 4: Performance comparison on reasoning and language-heavy benchmarks.
Method MMMU HellaS TCaps
Full Prec 50.44 77.63 152.00
GPTQ 32.22 51.29 55.06
GPTAQ 22.25 49.82 46.52
VLMQ 33.67 53.19 55.90
To assess performance on language-heavy and multimodal reasoning tasks, we evaluate VLMQ on HellaSwag [zellers2019hellaswag], TextCaps [zellers2019hellaswag], and MMMU [yue2023mmmu], respectively. These benchmarks cover pure language understanding, visually grounded language generation, and broad-domain multimodal reasoning, allowing us to evaluate the generalization ability of VLMQ beyond standard VQA-style tasks. These benchmarks cover pure language understanding, visually grounded language generation, and broad-domain multimodal reasoning, allowing us to evaluate the generalization ability of VLMQ beyond standard VQA-style tasks.
5.3 Ablation Study
Table 5: Ablation studies on importance-aware strategies.
Method LI Ratio DW Factor TextVQA DocVQA
Full Prec - - 82.02 93.89
GPTQ - - 71.93 74.90
GPTAQ - - 67.98 72.57
VLMQ $_{\text{naive}}$ 0.25 0.01 70.66 72.12
VLMQ $_{\text{naive}}$ 0.05 72.12 72.63
VLMQ $_{\text{naive}}$ 0.10 70.89 72.10
VLMQ $_{\text{naive}}$ 0.50 0.01 72.38 72.76
VLMQ $_{\text{naive}}$ 0.05 69.83 71.93
VLMQ $_{\text{naive}}$ 0.10 71.79 73.67
VLMQ $_{\text{naive}}$ 0.75 0.01 69.74 69.44
VLMQ $_{\text{naive}}$ 0.05 68.17 71.05
VLMQ $_{\text{naive}}$ 0.10 71.21 70.33
VLMQ (Ours) Grad-driven Grad-driven 74.80 75.76
5.3.1 Ablation on importance-aware strategies.
[…]
5.3.2 Ablation on importance factor type.
Table 6: Ablation study on the importance factor type. Reported results are under the INT-3 setting.
Factor Source Style Avg Acc $(\uparrow)$
Qwen2-VL-7B-Instruct, Avg Acc: 75.82%
Constant $\operatorname{Diag}({\mathbf{1}})$ 69.03%
Attention score FastV [chen2024fastv] 68.29%
Attention score PACT [dhouib2025pact] 67.07%
Gradient VLMQ (ours) 69.70%
As discussed in Section 4, in addition to gradient information, prior works such as FastV [chen2024fastv] and PACT [dhouib2025pact] incorporate attention score-based factors to estimate token significance. We evaluate the quantization performance with different factor types under INT-3 quantization in Table 6. For a fair comparison, all other settings are kept consistent across experiments. The results demonstrate that our formally proved gradient-driven importance factor preserves the performance of quantized models, while empirical attention score-based factors lead to substantial performance degradation. Extended results can be found in Table 14.
5.3.3 Ablation on backpropagation granularity.
To acquire the raw gradient, we have three potential target losses: 1) layer-wise distillation loss, 2) block-wise distillation loss, and 3) network-wise SFT loss (cross-entropy between predictions and ground truth labels). We provide the corresponding comparison in Table 7. Apparently, the block-wise manner, which is exactly used in the proposed VLMQ, provides an effective and efficient solution.
Table 7: Ablation study on backward granularity. Reported results are under the INT-3 setting.
Loss Type Avg Acc $(\uparrow)$ GPU Hours
Qwen2-VL-7B-Instruct, Avg Acc: 75.82%
${\mathcal{L}}_{\text{Layer}}=\operatorname{MSE}(\cdot)$ 67.77% 0.29
${\mathcal{L}}_{\text{Network}}=\operatorname{CE}(\cdot)$ 68.64% 0.70
${\mathcal{L}}_{\text{Block}}=\operatorname{MSE}(\cdot)$ 69.70% 0.21
We speculate the reasons behind the failure of layer-wise and network-wise manner lie in the lack of catching cross-layer dependency and overfitting in the limited calibration datasets. What’s more, our efficiency is guaranteed by the single lightweight block-wise backpropagation, which yields the lowest quantization hours among these three. Extended results are provided in Table 15.
5.3.4 Ablation on layer selection.
Table 8: Ablation study on layers to enhance under the INT-3 setting.
Breakpoint Layers Avg Acc $(\uparrow)$
Qwen2-VL-7B-Instruct, Avg Acc: 75.82%
- $\varnothing$ 69.03%
o_proj_in $\{\texttt{x\_proj}\}_{\texttt{x=q/k/v}}$ 55.54%
attn_out $\{\texttt{x\_proj}\}_{\texttt{x=q/k/v}}$ 56.21%
attn_out $\{\texttt{x\_proj}\}_{\texttt{x=q/k/v/o}}$ 69.70%
We also investigate the layers to be enhanced by our proposed VLMQ. As shown in Table 8, we find that placing the breakpoint at o_proj_in and skipping the enhancement of o_proj leads to failure. We speculate that breaking the integrity of a residual stream will severely damage performance. However, our block-split strategy still originates from an empirical setting like previous works [li2021brecq, li2024tesseraq, ding2023cbq]. A more fine-grained block-split strategy is required to be explored, which we leave to our future work. Extended results are provided in Table 16.
5.3.5 Ablation on precursor algorithms.
Table 9: Average accuracy improvements brought by VLMQ on INT2g128 quantized models.
Method Avg ($\uparrow$)
Qwen2-VL-7B-Instruct-INT2g128
GPTQ 55.43
GPTQ + VLMQ 57.76
GPTAQ 55.07
GPTAQ + VLMQ 56.51
Qwen2.5-VL-7B-Instruct-INT2g128
GPTQ 55.58
GPTQ + VLMQ 57.46
GPTAQ 41.69
GPTAQ + VLMQ 52.10
To better understand the role of precursor PTQ algorithms, we compare VLMQ when applied on top of GPTQ and GPTAQ under the INT2g128 setting (Table 9). GPTQ adopts layer-isolated calibration, whereas GPTAQ performs sequential layer-wise calibration using quantized inputs, which is beneficial at moderate bit-widths (e.g., INT3) but becomes unstable under ultra–low-bit settings due to amplified accumulated error. As a result, GPTAQ exhibits degraded performance at INT2, while GPTQ remains more reliable. Across both 7B and 7B-2.5 variants, VLMQ consistently improves each precursor by up to 10% on GPTAQ, demonstrating that our importance-aware mechanism is orthogonal to the calibration pipeline and enhances both symmetric and asymmetric PTQ formulations.
5.4 Efficiency
5.4.1 Quantization efficiency.
Table 10: Memory usage and quantization latency of different model sizes on a single H100 80GB GPU.
Size Metric GPU Cost